OpenStudy (anonymous):

The sum of two numbers is 18. the sum of their square is four more than 16 times the larger number. What is the number?

OpenStudy (anonymous):

How do I solve this using only one variable?

OpenStudy (anonymous):

Let the greater number be x Let he smaller number be y Needd to use 2 variables

OpenStudy (anonymous):

Form an equation of this first The sum of two numbers is 18.

OpenStudy (ikram002p):

form two equations asume numbers r x & y x+y =18 ..... assume x is the larger , y=18-x second equation x^2+y^2+4=16x

OpenStudy (anonymous):

Then what would i do next? transpose 16x?

OpenStudy (ikram002p):

so now substitute y value in the second equatoin \[x^2+(18-x)^2+4=16x\]

OpenStudy (ikram002p):

analyzed between brackets \[x^2+(324+x^2-36x)+4=16x\] transpose 16x \[2x^2 -52x+328=0\]

OpenStudy (ikram002p):

solve for x in the finall equation .

OpenStudy (anonymous):

x1= 10 and x2=16

OpenStudy (anonymous):

@ikram002p is this correct?

OpenStudy (ikram002p):

but sry i made a mistake the second equation is x^2+(18-x)^2-4= 16 x so the equation would be x^2 - 26x+160=0 So ur answer is correct good job.

myininaya (myininaya):

But the sum of those numbers he found he not 18.

myininaya (myininaya):

you also need to find y

myininaya (myininaya):

guess the number they are looking for is the larger number anyways

OpenStudy (ikram002p):

yeb he found 2 values for x now find 2 values for y x+y=18

OpenStudy (anonymous):

Ok. Thanks for the help!

OpenStudy (ikram002p):

ur wlc :) u such a smart one

OpenStudy (anonymous):

thanks